# Beyond Poisson: First-Passage Asymptotics of Renewal Shot Noise

### v1

@@ Abstract
The first-passage time (FPT) of a stochastic signal to a threshold is a fundamental observable across physics, biology, and finance. While renewal shot noise is a canonical model for such signals, analytical results for its FPT have remained confined to the Poisson (Markovian) case, despite the prevalence of non-Poisson arrival statistics in applications from neuronal spiking to gene expression. We break this long-standing barrier by deriving the first universal asymptotic formula for the mean FPT $\langle T_{b}\rangle$ to reach level $b$ for renewal shot noise with general arrival statistics and exponential marks. Our central result is a closed-form expression that reveals precisely how general inter-arrival statistics impact the naive Arrhenius law. We show that the short-time behavior of the interarrival distribution dictates universal scaling corrections, ranging from stretched-exponential to algebraic, that can dramatically accelerate threshold crossing. Furthermore, we argue and confirm numerically that the full FPT distribution becomes exponential at large thresholds, implying that $\langle T_{b}\rangle$ provides a complete asymptotic characterization. Our work, enabled by a novel exact solution for the moments of the noise, establishes a general framework for analyzing extreme events in non-Markovian systems with relaxation.
@@
$X(t)=\sum_{t_{i}\leq t}x_{i}\,e^{-\gamma(t-t_{i})},$ (1)
where $x_{i}$ are i.i.d. marks and interarrival times $\tau_{i}=t_{i+1}-t_{i}$ are i.i.d. with density $w(\tau)$ (see FIG. 1). This model embodies two essential features: impulsive bursts at random times, and relaxation between events. The classical Poisson case ($w(t)=re^{-rt}$) renders $X(t)$ Markovian, and its FPT statistics are well understood [17, 18, 19, 20, 21]. However, many applications exhibit strongly non-Poissonian arrival statistics, such as refractory periods in neuronal spiking or bursty transcription in gene expression [22, 23, 24, 25, 26, 11], which render the process genuinely non-Markovian. Despite decades of study, analytical progress on FPT statistics has remained confined to the Poisson case, with non-Poisson shot noise presenting a long-standing challenge, as is often the case for non-Markovian processes [27, 28, 29, 30]. While general, exact integral equations satisfied by the MFPT are known [21], these have proven to be intractable beyond the Poisson case.
Figure 1: A typical realization of renewal shot noise $X(t)$ (solid line). The vertical green line signals the FPT $T_{b}$, where $X(t)$ exceeds the threshold $b=5$ (horizontal red line) for the first time.
In this Letter, we break this barrier by deriving the first exact asymptotic expression for the mean FPT $\langle T_{b}\rangle$ for general renewal arrivals and exponential marks. Our result, given by a compact product formula Eq. (2), reveals explicitly how general interarrival statistics impact the FPT scaling, and in particular reveals universal deviations from the naive Arrhenius law. This breakthrough is enabled by a novel closed-form expression for the Laplace transform of all moments of $X(t)$, a result of independent interest that provides a powerful analytical tool for studying renewal shot noise.
Main result. We now present our main result for the MFPT $\langle T_{b}\rangle$. The marks are taken to be exponentially distributed, $x_{i}\sim\mathrm{Exp}(\lambda^{-1})$, and the arrival process has a finite mean rate $r\equiv\left(\int_{0}^{\infty}t\,w(t)\,dt\right)^{-1}$. The process may start at any value $0\leq X(0)\ll b$. For a function $f(t)$, we denote its Laplace transform by $\hat{f}(s)=\int_{0}^{\infty}e^{-st}f(t)\,dt$. Our central result is the following simple, exact asymptotic expression for the MFPT at large thresholds  [^1] The product term $\prod_{m=1}^{\lambda b}\!\bigl[1-\hat{w}(m\gamma)\bigr]$ is defined via analytic continuation when $\lambda b\notin\mathbb{N}$.:
@@
Figure 2: MFPT $\langle T_{b}\rangle$ of renewal shot noise with exponential marks (symbols: simulations; dashed lines: exact asymptotics, Eq. (2)). Interarrival times follow a Gamma distribution $w(t)=\tfrac{rk}{\Gamma(k)}(rkt)^{k-1}e^{-rkt}$ with mean rate $r$ and shape $k$, while marks are exponential with $\lambda=1$. Cases with $k=3$ (blue, $\gamma=1,r=0.4$) and $k=2$ (yellow, $\gamma=1.5,r=0.6$) highlight refractory effects ($w(0)=0$), whereas $k=0.75$ (green, $\gamma=4.5,r=1.6$) and the Poisson limit $k=1$ (red, $\gamma=2.5,r=1.2$) illustrate bursty dynamics. Statistical errors are smaller than the symbol size. The inset shows the simulated/theoretical ratio, confirming good quantitative agreement even at moderate $\lambda b$, with convergence speed set by $k$: larger $k$ suppresses short interarrivals and accelerates approach to the asymptotics.
@@
Second, in the limit of instantaneous relaxation ($\gamma\to\infty$), we have $\hat{w}(m\gamma)\to 0$, yielding the pure Arrhenius law $r\langle T_{b}\rangle\sim e^{\lambda b}=1/\mathbb{P}(x>b)$. Indeed, in this limit, each impulse is an upcrossing with probability $\mathbb{P}(x>b)=e^{-\lambda b}$. (ii) Equation (2) extends MFPT asymptotics for shot noise far beyond the Poisson case, providing (to our knowledge) the first closed analytic form valid for general renewal arrivals and exponential marks. Comparable asymptotics are extremely rare for non-Markovian processes [30, 29, 27, 32]. (iii) Although the full distribution of $T_{b}$ for arbitrary $b$ remains an open question for non-Markovian shot noise, our results provide a complete asymptotic characterization for large values of $b$, where we argue that crossings become asymptotically independent, implying that $T_{b}$ follows an exponential distribution. While this has been rigorously established only for Poisson arrivals [17, 18], it is a common phenomenon in rare-event limits [33, 34]. Our simulations FIG. 3 confirm that for large $b$,
@@
where $\langle T_{b}\rangle$ is now fully explicit (2). Thus, under the physically sound hypothesis that upcrossings become independent, our result (2) fully quantifies FPT statistics of shot noise with exponential marks in the large-threshold limit.
Figure 3: Numerical confirmation of the exponential distribution of $T_{b}$ for large $b$. The cumulative distribution function $\mathbb{P}(T_{b}<t)$ from simulations (in blue) is shown for a threshold $b=8$, with Gamma-distributed interarrivals (shape $k=2$, rate $r=0.6$), decay $\gamma=1.5$, and mark rate $\lambda=1$. The black dotted line is the expected exponential distribution $1-\exp(-t/\langle T_{b}\rangle)$, with $\langle T_{b}\rangle$ given by Eq. (2). The excellent agreement strongly supports our assertion (4) that the FPT distribution is asymptotically exponential for large $b$.
(iv) A central question in first-passage phenomena is how the MFPT deviates from the simple Arrhenius law $\langle T_{b}\rangle\propto e^{\lambda b}$ [30, 35, 36]. Assume the interarrival density behaves for short times as $w(t)\sim c\,t^{\kappa-1}$ with $\kappa>1/2$ (the case $0<\kappa\leq 1/2$ yields more intricate, stretched-exponential expressions and is treated in the SM). Our expression (2) then yields the universal asymptotic scaling
@@
where the prefactor $B$ (given in the SM) is nonuniversal and depends on the full interarrival law. The scaling itself, however, is universal and dictated solely by the short-time behavior of $w(t)$ through the parameters $c$ and $\kappa$. Physically, when $w(0)=0$ ($\kappa>1$), short interarrival times are suppressed, akin to refractory periods in neuronal spiking. In this regime, exponential relaxation prevents the accumulation of impulses, and threshold crossings are dominated by a single, rare large mark of size $\mathcal{O}(b)$, resulting in pure Arrhenius scaling. Conversely, when short gaps are prevalent ($\kappa\leq 1$), as observed in bursty gene transcription, crossings are driven by bursts of arrivals rather than isolated events. This cooperative mechanism leads to significant reductions in the mean first-passage time relative to the Arrhenius form: algebraic corrections when $\kappa=1$, and stretched-exponential corrections when $\kappa<1$.
@@
The structure of Eq. (Beyond Poisson: First-Passage Asymptotics of Renewal Shot Noise) reflects the underlying physics: it sums over all ways to distribute $n$ impulses among $k$ distinct arrivals ($1\leq k\leq n$). Each partition $n_{1}+\dots+n_{k}=n$ corresponds to a specific clustering pattern, with the $\hat{\psi}-$factors encoding the temporal structure.
@@
$\langle T_{b}\rangle\;\mathrel{\mathop{\kern 0.0pt\sim}\limits_{b\to\infty}}\;\frac{1}{r\,p(b)},$ (8)
where $p(b)$ is the probability that a single impulse in the stationary state pushes the process above the threshold $b$. Equation (8) rests on three key points. (i) For large $b$, crossings are rare and thus take a long time, so the process is near stationarity when a crossing occurs. (ii) Multiple crossings are exponentially less likely than a single crossing in the large-$b$ limit, so a crossing of $b$ at some time $t$ is very likely to be the first one. (iii) In the stationary regime, all impulses are equivalent, so the mean number of crossings per unit time is rp(b)r\,p(b). The quantity of interest is now
@@
The final step relies on an asymptotic duality (derived in SM) between truncated moment sums and integrals: for a random variable $Y$ with mgf finite for $t<\lambda$,
@@
the sum in (12) collapses, finally yielding Eq. (2).
Physical interpretation. We emphasize that the product form in Eq. (2) is not a mere mathematical artifact but reflects a clear physical mechanism, presented below. For large thresholds $b$, the MFPT is given by $\langle T_{b}\rangle\sim 1/(r\,p(b))$. To understand the product structure physically, we analyze how the crossing probability changes when we increment the threshold by one mark unit: $b^{+}\equiv b+1/\lambda$. The ratio $p(b^{+})/p(b)$ represents the conditional probability to cross $b^{+}$ given that $b$ has been crossed, and is governed by two distinct scenarios. Consider the overshoot $x_{0}$ remaining immediately after crossing $b$. Crucially, for exponential marks, $x_{0}$ is also exponentially distributed due to the memoryless property. This leads to two possible mechanisms: (S1) With probability $e^{-1}$, we have $x_{0}>1/\lambda$, meaning the same impulse that crossed $b$ also suffices to cross $b^{+}$. This contributes a term $e^{-1}p(b)$ and is responsible for the Arrhenius factor $e^{\lambda b}$. (S2) If instead $x_{0}\leq 1/\lambda$, the process starts below $b^{+}$ after crossing $b$. Since $X(t)$ rarely sits near such high values, crossing $b^{+}$ typically occurs through a burst of additional impulses before significant relaxation below $b$ can occur. This burst mechanism explains deviations from pure Arrhenius scaling. More precisely, in scenario (S2), crossing $b^{+}$ is achieved by $n\geq 1$ additional impulses with amplitudes $x_{1},\dots,x_{n}$ arriving at times $t_{1},\dots,t_{n}$. The interarrival times $\tau_{i}=t_{i}-t_{i-1}$ are most probably smaller than $(\lambda b^{+}\gamma)^{-1}$, the time needed for $X(t)$ starting from $b^{+}$ to relax by one mark unit $1/\lambda$. Since $\gamma\tau_{i}\ll 1$ for these relevant timescales, the condition for crossing $b^{+}$ after $n$ additional impulses becomes:
@@
$\mathbb{P}(A_{n})=e^{-\lambda\gamma b^{+}\sum_{k=1}^{n}\tau_{k}-1},$
which yields a term $\hat{w}(\lambda b^{+}\gamma)^{n}$ in (15). Combining both scenarios yields the recursion relation:
@@
which is exactly equivalent to our main result in Eq. (2)  [^2] Because only interarrival times with $\gamma\tau_{i}\ll 1$ matter, the MFPT expression (2) depends only on the local decay slope after an impulse, not on the full exponential form. The same asymptotics therefore hold in the large-$b$ limit if exponential relaxation is replaced by any smooth relaxation with the same initial slope. See [48] for the case of a constant slope.. This physical picture also clarifies why extending the explicit MFPT beyond the exponential–mark case solved here appears out of reach: for general marks the overshoot law is unknown and $\mathbb{P}(A_{n})$ lacks a closed form.

### v2

@@ Abstract
The first-passage time (FPT) of a stochastic signal to a threshold is a fundamental observable across physics, biology, and finance. While renewal shot noise is a canonical model for such signals, analytical results for its FPT have remained confined to the Poisson (Markovian) case, even though non-Poisson arrival statistics are common in systems from neuronal spiking to gene expression. Here, we overcome this long-standing limitation by deriving a universal asymptotic formula for the mean FPT (MFPT) $\langle T_{b}\rangle$ to reach level $b$ for renewal shot noise with arbitrary arrival statistics and exponential marks. Our central result is a simple, closed-form expression that exposes the physical mechanism by which temporal correlations in arrivals modulate the baseline Arrhenius law. We show that bursty arrivals introduce universal scaling corrections that markedly accelerate threshold crossings. In turn, non-bursty arrivals remain Arrhenius-like, directly linking temporal burstiness to Arrhenius scaling. Furthermore, we show and confirm numerically that the full FPT distribution becomes exponential at large thresholds, implying that $\langle T_{b}\rangle$ provides a complete asymptotic characterization. Our work, enabled by a novel exact expression for the moments of the noise, establishes a general framework for analyzing extreme events in non-Markovian systems with relaxation.
@@
$X(t)=\sum_{t_{i}\leq t}x_{i}\,e^{-\gamma(t-t_{i})},$
where $x_{i}$ are i.i.d. marks and interarrival times $\tau_{i}=t_{i+1}-t_{i}$ are i.i.d. with density $w(\tau)$ (see FIG. 1). This model embodies two essential features: impulsive bursts at random times, and relaxation between events. The classical Poisson case ($w(\tau)=re^{-r\tau}$) renders $X(t)$ Markovian, and its FPT statistics are well understood [17, 18, 19, 20, 21]. However, many applications exhibit strongly non-Poissonian arrival statistics, such as refractory periods in neuronal spiking or bursty transcription in gene expression [22, 23, 24, 25, 26, 11], which render the process genuinely non-Markovian. Despite decades of study, analytical progress on FPT statistics has remained confined to the Poisson case, with non-Poisson shot noise presenting a long-standing challenge, as is often the case for non-Markovian processes [27, 28, 29, 30]. While general, exact integral equations satisfied by the MFPT are known [21], these have proven to be intractable beyond the Poisson case: subsequent studies [31, 32, 33, 34] did not extract closed-form asymptotic behavior from them.
Figure 1: A typical realization of renewal shot noise $X(t)$, showing exponential relaxation between random impulsive events. The vertical green line signals the FPT $T_{b}$, where $X(t)$ exceeds the threshold $b=5$ (horizontal red line) for the first time.
In this Letter, we derive, to our knowledge, the first exact asymptotic expression for the MFPT $\langle T_{b}\rangle$ valid for general renewal arrivals and exponential marks. Our result, given by a compact product formula Eq. (1), reveals explicitly how general interarrival statistics impact the FPT scaling, and in particular reveals universal deviations from the baseline Arrhenius law. This breakthrough is enabled by a novel closed-form expression for the Laplace transform of all moments of $X(t)$, a result of independent interest that provides a powerful analytical tool for studying renewal shot noise.
Main result. We now present our main result for the MFPT $\langle T_{b}\rangle$. The marks are taken to be exponentially distributed, $x_{i}\sim\mathrm{Exp}(\lambda^{-1})$, and the arrival process has a finite mean rate $r\equiv\left(\int_{0}^{\infty}t\,w(t)\,dt\right)^{-1}$. The process may start at any value $0\leq X(0)\ll b$. For a function $f(t)$, we denote its Laplace transform by $\hat{f}(s)=\int_{0}^{\infty}e^{-st}f(t)\,dt$. Our central result is the following simple, exact asymptotic expression for the MFPT at large thresholds  [^1] We write $A(x)\sim B(x)$ for $x\to\infty$ if the ratio $A(x)/B(x)\to 1$. The product term $\prod_{m=1}^{\lambda b}\!\bigl[1-\hat{w}(m\gamma)\bigr]$ is defined via analytic continuation when $\lambda b\notin\mathbb{N}$. See SM for more details and an explicit analytical continuation in the case of Gamma distributed interarrivals considered on FIGs 2 and 3.:
@@
Figure 2: MFPT $\langle T_{b}\rangle$ of renewal shot noise with exponential marks (symbols: simulations; dashed lines: exact asymptotics, Eq. (1)). Interarrival times follow a Gamma distribution $w(t)=\tfrac{rk}{\Gamma(k)}(rkt)^{k-1}e^{-rkt}$ with shape parameter $k$ and mean rate chosen as $r=1.2/k$, while marks are exponential with $\lambda=1$. The blue and yellow cases highlight refractory effects ($w(0)=0$), whereas the green curve and the red Poisson baseline illustrate bursty dynamics ($w(0)>0$ or infinite). Statistical errors are smaller than the symbol size. The inset shows the ratio $\langle T_{b}\rangle/\langle T_{b}\rangle_{\mathrm{RE}}$, where $\langle T_{b}\rangle_{\mathrm{RE}}=1/(rp(b))$ is the rare-event estimate introduced in Eq. (7) and given by Eq. (1). In agreement with the analysis presented in the SM, one observes $\langle T_{b}\rangle\geq\langle T_{b}\rangle_{\mathrm{RE}}$ at finite $b$. As shown in the inset, the convergence $\langle T_{b}\rangle/\langle T_{b}\rangle_{\mathrm{RE}}\to 1$ is slower in the bursty regime $\kappa\leq 1$, consistent with enhanced clustering of the shots.
@@
Second, in the limit of instantaneous relaxation ($\gamma\to\infty$), we have $\hat{w}(m\gamma)\to 0$, yielding the pure Arrhenius law $r\langle T_{b}\rangle\sim e^{\lambda b}=1/\mathbb{P}(x>b)$. Indeed, in this limit, each impulse is an upcrossing with probability $\mathbb{P}(x>b)=e^{-\lambda b}$. (ii) Equation (1) extends MFPT asymptotics for shot noise far beyond the Poisson case, providing (to our knowledge) the first closed analytic form valid for general renewal arrivals and exponential marks. Comparable asymptotics are extremely rare for non-Markovian processes [30, 29, 27, 36]. (iii) Although the full distribution of $T_{b}$ for arbitrary $b$ remains an open problem for non-Markovian shot noise, our results provide a complete asymptotic characterization in the large-threshold limit. Indeed, we show in the Supplementary Material (SM) [54] that for $b\to\infty$, the FPT $T_{b}$ becomes exponentially distributed:
@@
where $\langle T_{b}\rangle$ is now fully explicit (1). Physically, this reflects the fact that upcrossings of a high threshold become asymptotically independent events. This property is well established for Poisson shot noise [17, 18] and more generally arises in rare-event limits [37, 38]. Our numerical simulations in FIG. 3 provide clear confirmation of this asymptotic behavior. Thus, our result (1) fully quantifies FPT statistics of shot noise with exponential marks in the large-threshold limit.
Figure 3: Numerical confirmation of the exponential form of the FPT distribution at large thresholds. Shown is the cumulative distribution function $\mathbb{P}(T_{b}>t)$ obtained from simulations (symbols) for $b=8$ and $\lambda=1$. The color code corresponds to different Gamma distributions for the interarrival times and is the same as in FIG. 2. Dotted lines show the exponential distribution (3): black lines use the empirical mean $\langle T_{b}\rangle$ measured in simulations, while colored lines use the theoretical MFPT predicted by Eq. (1). The excellent agreement, in particular between the black lines and the simulation data, provides strong numerical confirmation of our result (3) that the distribution of $T_{b}$ becomes exponential for large $b$. The inset displays $1-\mathbb{P}(T_{b}>t)$ in the small-$t$ regime, confirming the exponential form already well before $T_{b}$ approaches its mean value $\langle T_{b}\rangle$, which, as discussed in the SM, emerges rapidly. In contrast, the convergence of $\langle T_{b}\rangle$ to the rare-event estimate Eq. (1) is slower in the bursty regime $\kappa\leq 1$.
(iv) As detailed in the Physical interpretation section, Eq. (1) exposes a universal mechanism governing the kinetics of rare threshold crossings. The MFPT retains an Arrhenius-like scale $\langle T_{b}\rangle\propto e^{\lambda b}$, but the product term in Eq. (1) quantifies how temporal correlations (here, in the arrival process) modulate this baseline exponential scale, which is a central question in first-passage kinetics [30, 39, 40]. Assume the interarrival density behaves for short times as $w(t)\sim c\,t^{\kappa-1}$ with $\kappa>1/2$ (for $0<\kappa\leq 1/2$, Arrhenius modulation persists but leads to more intricate corrections detailed in the SM). Our expression (1) then yields the universal asymptotic scaling
@@
where the prefactor $B=B[w(t),\gamma]$ (given in the SM) is nonuniversal and depends on the full interarrival law. The scaling itself, however, is universal and dictated solely by the short-time behavior of $w(t)$ through the parameters $c$ and $\kappa$. Physically, when $w(0)=0$ ($\kappa>1$), short interarrival times are suppressed, akin to refractory periods in neuronal spiking. In this regime, exponential relaxation prevents the accumulation of impulses, and threshold crossings are dominated by a single, rare large mark of size $\mathcal{O}(b)$, resulting in pure Arrhenius scaling. Conversely, when short gaps are prevalent ($\kappa\leq 1$, with the baseline Poisson case included in $\kappa=1$), as observed in e.g. bursty gene transcription [11, 12], crossings are driven by bursts of arrivals rather than isolated events. This cooperative mechanism leads to substantial acceleration of threshold crossings: algebraic for $\kappa=1$, and stretched-exponential for $\kappa<1$, thereby establishing a direct quantitative link between arrival burstiness and suppression of the Arrhenius timescale $e^{\lambda b}/r$. As shown in the SM, this correction strengthens as $\kappa$ decreases. In the representative case $\kappa=1/2$, relevant to transport-limited gene dynamics [11] through the return statistics of a one-dimensional random walk, the interarrival density behaves as $w(t)\sim ct^{-1/2}+d$. The MFPT then acquires an additional polynomial correction: $\langle T_{b}\rangle\sim Be^{\lambda b-2c\sqrt{\tfrac{\pi\lambda b}{\gamma}}}(\lambda b)^{-d/\gamma}$.
@@
$\langle T_{b}\rangle\mathrel{\mathop{\kern 0.0pt\sim}\limits_{b\to\infty}}\langle T_{b}\rangle_{\mathrm{RE}}\equiv\frac{1}{r\,p(b)},$ (7)
where $p(b)$ is the probability that a single impulse, sampled from the stationary pre-jump distribution, pushes the process above the threshold $b$. The convergence of the actual MFPT $\langle T_{b}\rangle$ to this rare-event estimate is demonstrated in the SM, where it is shown that the generic inequality $\langle T_{b}\rangle\geq\langle T_{b}\rangle_{\mathrm{RE}}$ holds at finite $b$ (see FIGs. 2 and 3). We provide a physical interpretation of Eq. (7) here: it rests on three key points. (i) For large $b$, threshold crossings are rare and occur on timescales much larger than the time needed for the process to reach stationarity (which is independent of $b$). Thus, crossings take place in the stationary regime. (ii) In this rare-event limit, multiple crossings occurring within a short time window are exponentially unlikely, implying that an upcrossing of $b$ at time $t$ is overwhelmingly likely to be the first one. (iii) In the stationary regime, all impulses are statistically equivalent, so that the mean number of threshold crossings per unit time is simply rp(b)r\,p(b). Although this rare-event estimate is formally reminiscent of Kac-type relations [51, 52], which relate return times to a state $x$ to the inverse of the stationary density $1/P_{s}(x)$, we stress that Eq. (7) is of a different nature, as it concerns first-passage events rather than returns to the threshold $b$. The quantity of interest is now
@@
The final step relies on an asymptotic duality (derived in SM) between truncated moment sums and integrals: for a nonnegative random variable $Y$ with mgf $\langle e^{tY}\rangle$ finite for $t<\lambda$,
@@
the sum in the LHS of (11) collapses, yielding Eq. (1).
Physical interpretation. We emphasize that the product form in Eq. (1) is not a mere mathematical artifact but reflects a clear physical mechanism, presented below. For large thresholds $b$, the MFPT is given by the rare-event estimate $\langle T_{b}\rangle\sim\langle T_{b}\rangle_{\mathrm{RE}}=1/(r\,p(b))$. To understand the product structure physically, we analyze how the crossing probability changes when we increment the threshold by one mark unit: $b^{+}\equiv b+1/\lambda$. The ratio $p(b^{+})/p(b)$ represents the conditional probability to cross $b^{+}$ given that $b$ has been crossed, and is governed by two distinct scenarios. Consider the overshoot $x_{0}$ remaining immediately after crossing $b$. Crucially, for exponential marks, $x_{0}$ is also exponentially distributed due to the memoryless property. This leads to two possible mechanisms: (S1) With probability $e^{-1}$, we have $x_{0}>1/\lambda$, meaning the same impulse that crossed $b$ also suffices to cross $b^{+}$. This contributes a term $e^{-1}p(b)$ and is responsible for the Arrhenius factor $e^{\lambda b}$. (S2) If instead $x_{0}\leq 1/\lambda$, the process starts below $b^{+}$ after crossing $b$. Since $X(t)$ rarely sits near such high values, crossing $b^{+}$ typically occurs through a burst of additional impulses before significant relaxation below $b$ can occur. This burst mechanism explains deviations from pure Arrhenius scaling. More precisely, in scenario (S2), crossing $b^{+}$ is achieved by $n\geq 1$ additional impulses with amplitudes $x_{1},\dots,x_{n}$ arriving at times $t_{1},\dots,t_{n}$. The interarrival times $\tau_{i}=t_{i}-t_{i-1}$ are most probably smaller than $(\lambda b^{+}\gamma)^{-1}$, the time needed for $X(t)$ starting from $b^{+}$ to relax by one mark unit $1/\lambda$. Since $\gamma\tau_{i}\ll 1$ for these relevant timescales, the condition for crossing $b^{+}$ after exactly $n$ additional impulses becomes:
@@
$\mathbb{P}(A_{n})=e^{-\lambda\gamma b^{+}\left[\sum_{k=1}^{n}\tau_{k}\right]-1}.$ (15)
This allows the $n$-fold integral in (14) to decouple into a product of individual Laplace transforms, yielding a geometric sum of terms $e^{-1}\hat{w}(\lambda b^{+}\gamma)^{n}$ in (14). Combining both scenarios yields the recursion relation:
@@
which is exactly equivalent to our main result in Eq. (1). This physical picture enables two key insights. First, it clarifies why extending the explicit MFPT beyond the exponential-mark case solved here appears out of reach: for general marks, the overshoot law is unknown and $\mathbb{P}(A_{n})$ lacks a closed form. Second, because only interarrival times with $\gamma\tau_{i}\ll 1$ contribute in the large-$b$ limit, our exact result Eq. (1) depends solely on the local decay rate following each impulse, not on the full exponential form $x_{i}e^{-\gamma\tau_{i}}$. Consequently, Eq. (1) is robust to the precise form of the relaxation mechanism: it remains valid if the exponential relaxation is replaced by any smooth relaxation with the same initial slope $\gamma x_{i}$ (see [53] for the constant-slope case).
Conclusion. We have characterized the first-passage time $T_{b}$ to a high threshold $b$ for non-Markovian renewal shot noise with exponential marks. Our main results are: (i) an explicit asymptotic formula for the MFPT $\langle T_{b}\rangle$ (Eq. (1)), proven exact as $b\to\infty$ in the SM; (ii) the full FPT distribution becomes exponential for large $b$, i.e. $P(T_{b}>t)\sim e^{-t/\langle T_{b}\rangle}$ (Eq. (3)), confirmed analytically (see SM) and numerically (FIG. 3). By providing an explicit expression for the MFPT $\langle T_{b}\rangle$, our results fully characterize the distribution of $T_{b}$; (iii) exact closed-form expressions for all moments of renewal shot noise (Eqs. (Beyond Poisson: First-Passage Asymptotics of Renewal Shot Noise)–(6)), previously unavailable in the literature. These results expose how temporal correlations, that is, burstiness or refractoriness in arrivals, universally modulate the baseline Arrhenius scaling (Eq. (4)), establishing a direct link between microscopic arrival statistics and macroscopic extreme-event kinetics. Our framework enables quantitative predictions of rare threshold crossings in diverse non-Markovian systems, from neuronal spiking to gene expression bursts. A natural next step is quantifying finite-threshold corrections beyond the rare-event estimate $\langle T_{b}\rangle_{\mathrm{RE}}$, particularly in the bursty regime where such corrections could be of importance in quantifying phenotypic switching in transport-limited gene dynamics [11, 12].
Details and source code for the numerical simulations used to generate FIGs. 2 and 3 are available at https://github.com/julien-bremont/Shot-noise.